Re: 485 No Trace - some progress, question


Robert Simpson
 

Quick test: I tried IC U1560 in the working scope and got a trace, so it looks like some input is keeping it in blanking mode. On to more circuit analysis and tracing.
Bob

--- In TekScopes@yahoogroups.com, David wrote:

The Tek-Made Integrated Circuits Catalog has details and an internal
schematic for the 155-0012-00 Z-Axis Logic IC.

Pins 5 and 15 are high impedance logic inputs so the slightly higher
voltages you measured should not damage a replacement IC but before
trying a swap, I would check the circuits that generate the voltages
going into pins 5 and 15 although I am not sure what I would look for.
There is no reason for those pins to be high.

I would also check the area around U1585.

I do not see any evidence that U1560 is bad. I suspect an earlier
problem is keeping the CRT beam blanked.

On Wed, 30 Jan 2013 19:37:58 -0000, "Robert"
wrote:

Yes,
With further review of the schematic, I realized I was going into the focus control. Looking back, I see pin 8 of U1560 affects the Z out test point. Testing Pin 8 on the working scope varies about one volt. Pin 8 on the non working scope doesn't vary at all. I re-socketed this IC and the surrounding transistors. No effect.
Measurements for U1560:

U1560
Pin Working Non Working
1 -62mv +26mv
2 +1.1V +1.1V
3 +0.7V +0.3V
4 +51mv +48mv
5 +4.9mv +1.06V
6 +0.78V +0.79V
7 +0.96V +1.0V
8 +13.9V +14V
9 +3.6mv -1.0mv
10 +4.25V +4.2V
11 +14.9V +15V
12 +5.0V +4.9V
13 +4.9V +4.9V
14 +98mv +99mv
15 -187mv +928mv
16 -138mv +282mv

Question, given the above measurements, would it be safe to swap in the working U1550 without risk of damage?

Bob


--- In TekScopes@yahoogroups.com, David wrote:

I do not envy you trying to work through that z-axis amplifier. Having
a working 485 for comparison will help a lot.

R1703 is part of the focus circuit path. Z In (TP1741) and Z Out
(TP1781) are the input and output of the z-axis amplifier.

If the z-axis amplifier transistors are socketted, I would pull and
check them individually for shorts and opens using the diode test
function on my multimeter.

On Wed, 30 Jan 2013 16:27:26 -0000, "Robert"
wrote:

Last night tried the Grid Bias adjustment procedure item 10 on page 5-10 and found a problem. The non working scope has no change at the Z Out DC test point when adjusting the A channel intensity. My working 485 changes as expected when adjusting the front panel A channel intensity control. The non working scope stays the same at about 8.2 volts. 8.2 volts is about the same as the working scope with the front panel A channel intensity all the way down. So the beam must be in cutoff condition.

So far have traced to R1703 and the voltage varies OK to that point (same as the working 485). I will continue tracing today.
Bob

Join TekScopes@groups.io to automatically receive all group messages.